OpenStudy (anonymous):

Please help... Find the slope of the tangent line to the parabola y= 4x-x^2 at the point (1,3)

OpenStudy (anonymous):

You want the derivative :\[\frac{dy}{dx}=\frac{d}{dx}(4x - x^{2})=4-2x\]let x = 1\[\frac{dy}{dx}=4-2\times1 = 2\]which is the slope of the tangent line at (1,3)